# NeuRAD: Neural Rendering for Autonomous Driving

### v2

@@ 4.4 Ablations
We validate the effectiveness of some key components in Tab. 4. To avoid biases toward any specific dataset, we report averaged metrics from sequences from all five datasets considered in this work. We select 4 diverse sequences from each dataset, see details in Appendix B. Our full model corresponds to the model used in all prior experiments and strikes a good balance between run-time and performance. We see that the CNN decoder (a) significantly increases both quality and speed, by requiring significantly fewer rays and allowing for interaction between rays. Accurate sensor modeling is also very important, as each of our contributions in that area provide complementary performance boost: considering rolling shutter (b), modeling each ray as a frustum (c), per-sensor appearance embeddings (d), and considering lidar rays that did not return (e). We also demonstrate that replacing individual actor hash grids with a single 4D hash grid (f) has no detrimental impact on quality, while significantly increasing training speed. Finally, we replace our SDF with a NeRF-like density formulation (g). The performance is overall almost identical and shows that our model can be configured to either of these field representations depending on the need. If we desire to extract surfaces from our model, we can use an SDF, but if our scenes are dominated by fog, transparent surfaces, or other effects where an SDF breaks down, we can fall back to a density formulation. Interestingly, our ablations only show a modest impact of considering rolling shutter. However, the effect is clearly evident at closer inspection of the qualitative results shown in Fig. 3. Here it is apparent that both the renderings and underlying geometry break down without considering this effect. More results, as well as failure cases, can be found in Appendix F.
@@ 5 Conclusions
In this paper, we have proposed NeuRAD, a neural simulator tailored specifically for dynamic autonomous driving (AD) data. The model jointly handles lidar and camera data in ${}^{\circ}$ and decomposes the world into its static and dynamic elements, allowing the creation of sensor-realistic editable clones of real world driving scenarios. NeuRAD incorporates novel modeling of various sensor phenomena including beam divergence, ray dropping, and rolling shutters, all increasing the quality during novel view synthesis. We demonstrate NeuRAD’s efficacy and robustness by obtaining state-of-the-art performance on five publicly AD datasets, using a single set of hyperparameters. Lastly, we publicly release our source-code to foster more research into NeRFs for AD.
@@
Acknowledegments
We thank Maryam Fatemi for the valuable feedback. Further, this work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at NSC Berzelius, partially funded by the Swedish Research Council, grant agreement no. 2022-06725.
@@
In the supplementary material, we provide implementation details for our method and baselines, evaluation details, and additional results. In Appendix A, we describe our network architecture more closely and provide hyperparameter values. In Appendix B, we provide details on the experimental setting. Then, in Appendix C, we provide details on our baseline implementation. We closely describe the process of inferring lidar rays that did not return in Appendix D. Next, we cover additional details of our proposed rolling shutter modeling in Appendix E. Last, in Appendix F, we showcase additional results and highlight some limitations of our method.
@@ Appendix B Evaluation details
Pandaset (UniSim): UniSim uses a simple evaluation protocol, where the entire sequence is used, with every other frame selected for evaluation and the remaining half of the frames for training. The authors report numbers for the front camera and the ${}^{\circ}$ lidar on the following sequences: 001, 011, 016, 028, 053, 063, 084, 106, 123, 158. We call this protocol Panda FC, and additionally report Panda 360 results, with all 6 cameras (and the ${}^{\circ}$ lidar). For the backward-facing camera, we crop away 250 pixels from the bottom of the image, as this mainly shows the trunk of the ego vehicle.
@@
Ablation Dataset: We perform all ablations on 20 sequences, four from each dataset considered above. We use sequences 001, 011, 063, 106 for PandaSet, 0164, 0209, 0359, 0916 for nuScenes, 0006, 0010, $0$, 0002 for KITTI, 000030, 000221, 000657, 000005 for ZOD, and 280269f9-6111-311d-b351-ce9f63f88c81, 185d3943-dd15-397a-8b2e-69cd86628fb7, 05fa5048-f355-3274-b565-c0ddc547b315, 0b86f508-5df9-4a46-bc59-5b9536dbde9f for Argoverse 2. Here, we no longer adopt the dataset-specific evaluation protocols corresponding to each SoTA method. Instead, we evaluate on the full sequences, on all available sensors, alternating frames for training and evaluation. The exception is nuScenes, where we find the provided poses to be too poor to train on the full sequences. If we optimize poses during training, we get qualitatively good results, and strong FID scores, but poor reconstruction scores due to misalignment between the learned poses and the evaluation poses, see Appendix F for a more detailed exposition. Therefore, we re-use S-NeRF’s shortened evaluation protocol, where this problem is mostly avoided, and leave the problem of proper evaluation on nuScenes for future work.
@@ Appendix D Inferring ray drop
Diode index assigment: All datasets considered in this work use spinning lidars, where a set of diodes are rotated ${}^{\circ}$ around the sensor and each diode is mounted at a fixed elevation angle. Typically, all diodes (or channels) transmit the same number of beams each revolution, where the number depends on the sensors’ horizontal resolution. To use this information for inferring missing rays, we need to assign each return to its diode index. For most datasets considered here [5, 1, 43], this information is present in the raw data. However, for the other [45], we must predict diode assignment based on the points’ elevation. As there is no ground truth available for this information, we use qualitative inspections to verify the correctness of the procedures outlined below.
@@ Appendix E Modeling rolling shutter
Appendix F Additional results
@@
Proposal sampling: To efficiently allocate samples along each ray, we use two rounds of proposal sampling. For comparison, UniSim [47] instead samples along the rays uniformly and relies on a lidar-based occupancy grid to prune samples far from the detected surfaces. Although the occupancy grid is fast to evaluate, it has two shortcomings. First, the method struggles with surfaces far from any lidar points. In the case of UniSim, the RGB values must instead be captured by the sky field, effectively placing the geometry far away regardless of its true position. The upper row of Fig. 7 shows an example of this, where a utility pole becomes very blurry without proposal sampling. Second, uniform sampling is not well suited for recovering thin structures or fine details of close-up surfaces. Doing so would require drawing samples very densely, which, instead, scales poorly with computational requirements. We examine both failure cases in Fig. 7, with thin power lines in the upper row and close-ups of vehicles in the lower row.

### v3

@@ 4.4 Ablations
We validate the effectiveness of some key components in Tab. 4. To avoid biases toward any specific dataset, we report averaged metrics from sequences from all five datasets considered in this work. We select 4 diverse sequences from each dataset, see details in Appendix B. Our full model corresponds to the model used in all prior experiments and strikes a good balance between run-time and performance. We see that the CNN decoder (a) significantly increases both quality and speed, by requiring significantly fewer rays and allowing for interaction between rays. Accurate sensor modeling is also very important, as each of our contributions in that area provide complementary performance boost: considering rolling shutter (b) or lidar rays that did not return (e), modeling each ray as a frustum (c) and per-sensor appearance embeddings (d). We also demonstrate that replacing individual actor hash grids with a single 4D hash grid (f) has no detrimental impact on quality, while significantly increasing training speed. Finally, we replace our SDF with a NeRF-like density formulation (g). The performance is overall almost identical and shows that our model can be configured to either of these field representations depending on the need. If we desire to extract surfaces from our model, we can use an SDF, but if our scenes are dominated by fog, transparent surfaces, or other effects where an SDF breaks down, we can fall back to a density formulation. Interestingly, our ablations only show a modest impact of considering rolling shutter. However, upon closer inspection of the qualitative results, see Fig. 3, it is apparent that both the renderings and underlying geometry break down without considering this effect.
@@ 5 Conclusions
In this paper, we have proposed NeuRAD, a neural simulator tailored specifically for dynamic autonomous driving (AD) data. The model jointly handles lidar and camera data in 360 and decomposes the world into its static and dynamic elements, allowing the creation of sensor-realistic editable clones of real world driving scenarios. NeuRAD incorporates novel modeling of various sensor phenomena including beam divergence, ray dropping, and rolling shutters, all increasing the quality during novel view synthesis. We demonstrate NeuRAD’s efficacy and robustness by obtaining state-of-the-art performance on five publicly AD datasets, using a single set of hyperparameters. Lastly, we publicly release our source-code to foster more research into NeRFs for AD.
@@
Acknowledegments: We thank Maryam Fatemi for valuable feedback. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Computational resources were provided by NAISS at NSC Berzelius, partially funded by the Swedish Research Council, grant agreement no. 2022-06725.
@@
In the supplementary material, we provide implementation details for our method and baselines, evaluation details, and additional results. In Appendix A, we describe our network architecture more closely and provide hyperparameter values. In Appendix B, we provide details on the experimental setting. Then, in Appendix C, we provide details on our baseline implementation. We closely describe the process of inferring lidar rays that did not return in Appendix D. Next, we cover additional details of our proposed rolling shutter modeling in Appendix E. Last, in Appendix G, we showcase additional results and highlight some limitations of our method.
@@ Appendix A Implementation details
NeuRAD-2x: We upscale NeuRAD in a straightforward manner – by doubling the size of all hash tables, thereby approximately doubling the model’s parameter count. As this model is primarily intended for long sequences and large scenes, we also double the resolution of each level of the static hashgrid. To accommodate the expanded model complexity, we extend the training to 50,000 iterations and adjust the warm-up periods correspondingly. All other hyperparameters remain the same. We find that while further scaling offers benefits in some cases, it leads to diminishing returns in others.
@@ Appendix B Evaluation details
Pandaset (UniSim): UniSim uses a simple evaluation protocol, where the entire sequence is used, with every other frame selected for evaluation and the remaining half of the frames for training. The authors report numbers for the front camera and the 360 lidar on the following sequences: 001, 011, 016, 028, 053, 063, 084, 106, 123, 158. We call this protocol Panda FC, and additionally report Panda 360 results, with all 6 cameras (and the 360 lidar). For the backward-facing camera, we crop away 250 pixels from the bottom of the image, as this mainly shows the trunk of the ego vehicle.
@@
Ablation Dataset: We perform all ablations on 20 sequences, four from each dataset considered above. We use sequences 001, 011, 063, 106 for PandaSet, 0164, 0209, 0359, 0916 for nuScenes, 0006, 0010, $0$, 0002 for KITTI, 000030, 000221, 000657, 000005 for ZOD, and 280269f9-6111-311d-b351-ce9f63f88c81, 185d3943-dd15-397a-8b2e-69cd86628fb7, 05fa5048-f355-3274-b565-c0ddc547b315, 0b86f508-5df9-4a46-bc59-5b9536dbde9f for Argoverse 2. Here, we no longer adopt the dataset-specific evaluation protocols corresponding to each SoTA method. Instead, we evaluate on the full sequences, on all available sensors, alternating frames for training and evaluation. The exception is nuScenes, where we find the provided poses to be too poor to train on the full sequences. If we optimize poses during training, we get qualitatively good results, and strong FID scores, but poor reconstruction scores due to misalignment between the learned poses and the evaluation poses, see Appendix G for a more detailed exposition. Therefore, we re-use S-NeRF’s shortened evaluation protocol, where this problem is mostly avoided, and leave the problem of proper evaluation on nuScenes for future work.
@@ Appendix D Inferring ray drop
Diode index assigment: All datasets considered in this work use spinning lidars, where a set of diodes are rotated 360 around the sensor and each diode is mounted at a fixed elevation angle. Typically, all diodes (or channels) transmit the same number of beams each revolution, where the number depends on the sensors’ horizontal resolution. To use this information for inferring missing rays, we need to assign each return to its diode index. For most datasets considered here [5, 1, 43], this information is present in the raw data. However, for the other [45], we must predict diode assignment based on the points’ elevation. As there is no ground truth available for this information, we use qualitative inspections to verify the correctness of the procedures outlined below.
@@ Appendix E Modeling rolling shutter
Refer to captionRefer to captionRefer to caption
Figure 7: Qualitative comparison between NeuRAD and UniSim across three Pandaset sequences (016, 028, 158). Displayed are the front left, front center, and front right camera perspectives. NeuRAD overall captures more details than UniSim, although the difference is not dramatic for the front camera. However, as highlighted by red boxes, NeuRAD clearly outperforms UniSim for side cameras.
Appendix F Simulation gap
In the following, we show results for the simulation gap. To study the real2sim gap, we train the 3D object detector BEVFormer on real images from PandaSet-360 and evaluate its object detection performance on synthesized images from the ten sequences used for NVS (not part of the training set for BEVFormer). For BEVFormer, we use the official implementation [^2] https://github.com/fundamentalvision/BEVFormer and the small version of the model. In Tab. 6 we see that the detector achieves similar validation performance for the real images and the synthesized images from NeuRAD. UniSim, struggling in the 360 setting, exhibits a larger gap.
Further, for a more general and dataset-agnostic evaluation, we use a zero-shot depth estimator, DepthAnything (DA). We measure the agreement between depth estimations on synthesized and real images using the standard $\delta_{1}$ metric. Tab. 7 shows consistent depth estimations, indicating a low domain gap across several datasets. For reference, DA reports $\delta_{1}\!=\!0.947$ when comparing against ground truth depth. We find these studies to give valuable insights and will include them in the manuscript.
Table 6: Real2Sim gap: BEVFormer (mAP) on different images.
Data source NeuRAD UniSim Real
mAP 32.0 30.1 32.4
Table 7: Real2Sim gap: DepthAnything relative depth ($\delta_{1}\!\uparrow$).
PandaFC Panda360 AV2 ZOD NuScenes KITTI
UniSim 0.927 0.872 0.860 0.901 - -
Neurad 0.968 0.944 0.928 0.958 0.894 0.947
Appendix G Additional results
@@
Comparison with UniSim: We begin with a direct qualitative comparison between NeuRAD and UniSim [47] as depicted in Fig. 7. For the front camera, the distinction in quality is subtle but observable; NeuRAD demonstrates superior image clarity, exhibiting notably less noise and artifact presence. In contrast, the disparity in quality is more pronounced with the side cameras. Here, NeuRAD’s output markedly surpasses UniSim, which is particularly evident in the highlighted areas where UniSim exhibits significant motion blur and visual distortion that NeuRAD effectively mitigates.
Proposal sampling: To efficiently allocate samples along each ray, we use two rounds of proposal sampling. For comparison, UniSim instead samples along the rays uniformly and relies on a lidar-based occupancy grid to prune samples far from the detected surfaces. Although the occupancy grid is fast to evaluate, it has two shortcomings. First, the method struggles with surfaces far from any lidar points. In the case of UniSim, the RGB values must instead be captured by the sky field, effectively placing the geometry far away regardless of its true position. The upper row of Fig. 8 shows an example of this, where a utility pole becomes very blurry without proposal sampling. Second, uniform sampling is not well suited for recovering thin structures or fine details of close-up surfaces. Doing so would require drawing samples very densely, which, instead, scales poorly with computational requirements. We examine both failure cases in Fig. 8, with thin power lines in the upper row and close-ups of vehicles in the lower row.